OpenStudy (anonymous):

the measures of the angles of a triangle are in the extended ratio of 7 : 9 : 10. Find the measures of the angles of the triangle

OpenStudy (gudden):

using angle sum property it means 7x + 9x +10x = 180 solve it further to get the answer @jmartinez2010 ... if u have any problem.. i m there for u then..

OpenStudy (gudden):

come on give it a try...

OpenStudy (anonymous):

26x=180

OpenStudy (gudden):

now get the value of x

OpenStudy (gudden):

=> x = 180 /26 so 10x = 1800/26 = 69.23 9x = 180*9/26 = 62.30 7x = 180*7/26 = 48.46 so that's ur answer buddy..!